Find a basis and the dimension of the subspace w of V spanned by the matrices

A=[ 1 2
-1 3]
B=[ 2 5
1 -1]
C=[3 4
-2 5]
1
Expert's answer
2020-12-23T18:04:36-0500

Let us find a basis and the dimension of the subspace "W" of "V" spanned by the matrices:


"A=\left[\begin{array}{cc} 1 & 2 \\ -1 & 3 \end{array}\right]", "B=\left[\begin{array}{cc} 2 & 5 \\ 1 & -1 \end{array}\right]", "C=\left[\begin{array}{cc} 3 & 4 \\ -2 & 5 \end{array}\right]" .


Let us consider a linear combination "kA+mB+nC=0." It follows that


"k\left[\begin{array}{cc} 1 & 2 \\ -1 & 3 \end{array}\right]+m\left[\begin{array}{cc} 2 & 5 \\ 1 & -1 \end{array}\right]+n\left[\begin{array}{cc} 3 & 4 \\ -2 & 5 \end{array}\right]=\left[\begin{array}{cc} 0 & 0 \\ 0 & 0 \end{array}\right]"


"\left[\begin{array}{cc} k+2m+3n & 2k+5m+4n \\ -k+m-2n & 3k-m+5n \end{array}\right]

=\left[\begin{array}{cc} 0 & 0 \\ 0 & 0 \end{array}\right]"


Therefore, we have the following system:


"\begin{cases}k+2m+3n=0\\ 2k+5m+4n=0 \\ -k+m-2n=0 \\ 3k-m+5n=0 \end{cases}" which is equivalent to "\begin{cases}k+2m+3n=0\\ m-2n=0 \\ 3m+n=0 \\ -7m-4n=0 \end{cases}" and to


"\begin{cases}k+2m+3n=0\\ m-2n=0 \\ 7n=0 \\ -18n=0 \end{cases}"


It follows that "n=m=k=0". Consequently, the matrices are linearly independent, and thus form a basis of the subspace "W" of "V" spanned by them. Also "\dim W=3."